Green Garden Company is looking at potential supply chain partners located either at New Orleans, Houston, or Dallas. They have estimated the following parameters for each location. For what demand range is each option the best?

Location

Fixed Cost

Variable Cost

New Orleans

350,000

50

Houston

450,000

40

Dallas

250,000

70

  1. Step 1: calculate indifference points (3 points)
  2. Step 2: draw the total cost line for each location (3 points)
  3. Step 3: identify the demand range for each location (3 points)
Answer #1

Indifference point between Dallas and New orleans ( assuming that it is N)

costs are equal at Indifference point

250000+70N= 350000+50N

N = 100000/20 =5000 units

Indifference point between Houston and New orleans ( assuming that it is N)

450000+40N = 350000+50N

100000/10 =10000

Dallas is better for 0-4999 units

New Orleans is better for 5001-9999 units

Houston is better for 10001 units upwards